OpenStudy (anonymous):

y varies directly with x, and y = 2 when x = 5. What is the value of x when y = 15?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

y varies directly with x so this means y = kx where k is some fixed number

jimthompson5910 (jim_thompson5910):

we are told that y = 2 when x = 5 so y = kx 2 = k*5 solve for k

OpenStudy (anonymous):

so 6?

jimthompson5910 (jim_thompson5910):

when you solved for k, what did you get

OpenStudy (anonymous):

wait how do i solve for k

jimthompson5910 (jim_thompson5910):

2 = k*5 divide both sides by 5 to isolate k

jimthompson5910 (jim_thompson5910):

k = ???

OpenStudy (anonymous):

2.5?

jimthompson5910 (jim_thompson5910):

you divided in the wrong order

OpenStudy (anonymous):

0.4?

jimthompson5910 (jim_thompson5910):

good, so y = kx turns into y = 0.4x

jimthompson5910 (jim_thompson5910):

now you want to find x when y = 15, so y = 0.4x 15 = 0.4x now solve for x

OpenStudy (anonymous):

37.5?

jimthompson5910 (jim_thompson5910):

you nailed it you can also represent 37.5 as the fraction 75/2

jimthompson5910 (jim_thompson5910):

so x = 37.5 or x = 75/2

OpenStudy (anonymous):

ok thank you(:

jimthompson5910 (jim_thompson5910):

you're welcome